# Pauli Hamiltonians with an Aharonov-Bohm flux

### v1

@@ A.1.1. Linear transformations
Using the basic algebraic identities $\{\sigma_{h},\sigma_{m}\}=2\delta_{hm}\mathds{1}$ and $[\sigma_{h},\sigma_{m}]=2i\varepsilon_{hmr}\sigma_{r}$, from (A.7) we deduce $T_{hj}T_{mj}=\delta_{hm}$ and $S(\mathbf{x})\,\sigma_{p}\,{S}^{-1}(\mathbf{x})\,\varepsilon_{j\ell p}\,T_{hj}T_{m\ell}=\varepsilon_{hmr}\,\sigma_{r}$. These are equivalent to
@@ A.2. Symmetries of Dirac Hamiltonians
$e^{-i\bm{\eta}\cdot\bm{\sigma}}\bm{\sigma}\,e^{i\bm{\eta}\cdot\bm{\sigma}}=\cos(2\eta)\,\bm{\sigma}-\tfrac{1}{\eta}\sin(2\eta)\,\bm{\eta}\wedge\bm{\sigma}+\tfrac{1}{\eta^{2}}\big{(}1-\cos(2\eta)\big{)}\,\left(\bm{\eta}\cdot\bm{\sigma}\right)\bm{\eta}.$ (A.23)

### v3

@@ A.1.1. Linear transformations
Using the basic algebraic identities $\{\sigma_{h},\sigma_{m}\}=2\delta_{hm}\mathds{1}$ and $[\sigma_{h},\sigma_{m}]=2i\varepsilon_{hmr}\sigma_{r}$, from (A.7) we deduce $T_{hj}T_{mj}=\delta_{hm}$ and $S(\mathbf{x})\,\sigma_{\mathrm{P}}\,{S}^{-1}(\mathbf{x})\,\varepsilon_{j\ell p}\,T_{hj}T_{m\ell}=\varepsilon_{hmr}\,\sigma_{r}$. These are equivalent to
@@ A.2. Symmetries of Dirac Hamiltonians
$e^{-i\bm{\eta}\cdot\bm{\sigma}}\bm{\sigma}\,e^{i\bm{\eta}\cdot\bm{\sigma}}=\cos(2\eta)\,\bm{\sigma}-\tfrac{1}{\eta}\sin(2\eta)\,\bm{\eta}\land\bm{\sigma}+\tfrac{1}{\eta^{2}}\big{(}1-\cos(2\eta)\big{)}\,\left(\bm{\eta}\cdot\bm{\sigma}\right)\bm{\eta}.$ (A.23)